OpenStudy (gypsylovesyoda):

WELL MEDAL AND FAN. A larger number is double the sum of 3 and a smaller number. The larger number is 2 less than 3 times the smaller number. If y represents the larger number and x represents the smaller number, which equations model the situation? Check all that apply.

OpenStudy (anonymous):

lol no one want to help with this one huh?

OpenStudy (gypsylovesyoda):

nope

OpenStudy (anonymous):

ok lets do it

OpenStudy (anonymous):

"the sum of 3 and the smaller number" you are told that the smaller number is \(x\) right?

OpenStudy (gypsylovesyoda):

please stay the than im writing is long so please wat.

OpenStudy (anonymous):

ok i stay

OpenStudy (gypsylovesyoda):

a.y=3x-3 b.3x-y=2 c.3x-y=-2 d.y=2-3x e.y=2(x+3)

OpenStudy (anonymous):

ok the "sum of 3 and the smaller number" sum means add, i.e. + the smaller number is x, so \(x+3\) then two times that is \(2(x+3)\) so one equation is \[y=2(x+3)\]

OpenStudy (anonymous):

The larger number is 2 less than 3 times the smaller number.

OpenStudy (anonymous):

3 times the smaller number is written as \(3x\) two less than that means \(3x-2\) so another equation is \[y=3x-2\]

OpenStudy (anonymous):

but they are trying to trick you by writing \[y=3x-2\] as \[3x-y=2\] which is the same thing

OpenStudy (gypsylovesyoda):

oh ok i was going to say i dont have that one.

OpenStudy (gypsylovesyoda):

are there eney more.

OpenStudy (anonymous):

oh idk let me check

OpenStudy (anonymous):

no i don't see any

OpenStudy (gypsylovesyoda):

ok thank you

OpenStudy (gypsylovesyoda):

well you help me with another one if you can.

OpenStudy (anonymous):

sure why not?

OpenStudy (anonymous):

is it as annoying as that last one?

OpenStudy (gypsylovesyoda):

not sure tell me if its to hard

OpenStudy (gypsylovesyoda):

Twice a number added to a smaller number is 5. The difference of 5 times the smaller number and the larger number is 3. Let x represent the smaller number and y represent the larger number. Which equations represent the situation?

OpenStudy (anonymous):

ok now lets go slow so you can do it

OpenStudy (gypsylovesyoda):

ok

OpenStudy (anonymous):

twice a number lets call that number \(y\) how do you write "twice y"?

OpenStudy (gypsylovesyoda):

what do you mean

OpenStudy (anonymous):

how do you write "two times a number y " in math hint, "three times a number y" i would write as \(3y\)

OpenStudy (gypsylovesyoda):

ok

OpenStudy (anonymous):

ok so how to you write "twice a number" if you are calling the number "y"?

OpenStudy (gypsylovesyoda):

never mid i have to go because i only have 20min in my quiz and i steel have 6 questions left.

OpenStudy (gypsylovesyoda):

thanks for helping tho.